# HARec: Hyperbolic Graph-LLM Alignment for Exploration and Exploitation in Recommender Systems

### v1

@@ Abstract.
Modern recommendation systems often create information cocoons, limiting users’ exposure to diverse content. To enhance user experience, a crucial challenge is developing systems that can balance content exploration and exploitation, allowing users to adjust their recommendation preferences. Intuitively, this balance can be achieved through a tree-structured representation, where depth search facilitates exploitation and breadth search enables exploration. However, current works face two challenges to achieve this target: (1) Euclidean methods fail to fully capture hierarchical structures and lack flexibility in balancing exploration-exploitation, while (2) hyperbolic approaches, despite better hierarchical modeling, suffer from insufficient semantic alignment due to their reliance on Euclidean text encoders. To address these challenges, we propose HARec, a hyperbolic representation learning framework that jointly aligns user-item collaborative information with textual descriptions in hyperbolic space. Our framework introduces two key technique novelty: (1) a hierarchical-aware graph-llm alignment mechanism that enables better hierarchical representation, and (2) a hyperbolic hierarchical tree structure that facilitates user-adjustable exploration-exploitation trade-offs. Extensive experiments demonstrate that HARec consistently outperforms both Euclidean and hyperbolic baselines, achieving up to 5.49% improvement in utility metrics and 11.39% increase in diversity metrics.
Recommender Systems, Representation Learning, Hyperbolic Space
Authors affiliated with Snap Inc. served in advisory roles only for this work.
Corresponding author
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: ; Aug 3–Aug 7, 2025; Woodstock, NYccs: Information systems Recommender systemsccs: Information systems Collaborative filteringccs: Mathematics of computing Geometric topologyccs: Information systems Personalization
1. INTRODUCTION
@@
As online platforms grow exponentially, users are overwhelmed by an influx of information, resulting in information cocoons that restrict their exposure to diverse and novel content (Kunaver and Požrl, 2017; Castells et al., 2021).
@@
Traditional recommendation approaches have primarily relied on collaborative filtering and interaction data (He et al., 2020; Koren et al., 2009), where users with similar historical preferences are expected to share future interests (Chen et al., 2021). However, these methods face two critical challenges in improving user experience. First, they often fail to capture the rich semantic relationships between items and evolving user preferences. For example, a jazz enthusiast might wish to explore rock music, yet current models struggle to understand these semantic connections and adapt to such evolving preferences (as illustrated in Figure 1). Second, existing approaches inadequately capture the underlying hierarchical structure inherent in user-item networks (Zheng et al., 2017; Unger and Tuzhilin, 2020), a critical factor for improving performance and personalization.
The emergence of large language models (LLMs) offers a solution to the limitations of semantic understanding. However, integrating LLM-derived semantic features with traditional collaborative signals remains challenging, especially when capturing the hierarchical structure of user preferences and item relationships.
@@
Hyperbolic space has emerged as a promising solution to these issues. It is a powerful tool for modeling tree-like structures and power-law distributions (Krioukov et al., 2010; Sun et al., 2021; Yang et al., 2022b, a), where its exponential volume growth naturally aligns with hierarchical, scale-free structures (Ganea et al., 2018; Chami et al., 2019; Liu et al., 2019; Zhang et al., 2021; Yang et al., 2023). In hyperbolic space, distances increase exponentially from the origin, allowing popular items or exploratory users to be positioned closer to the origin, while niche items or users with focused interests spread toward the boundaries. This spatial configuration enables hyperbolic graph neural network (GNN) models to perceive hierarchical relationships effectively, enhancing recommendation performance (Sun et al., 2021; Yang et al., 2022a, b).
@@
However, current hyperbolic recommendation systems still face two key limitations. First, hyperbolic approaches often lack proper semantic alignment due to their dependence on Euclidean text encoders, which struggle to integrate within hyperbolic space. This mismatch leaves existing hyperbolic models vulnerable to noise in collaborative information, as they fail to capture semantic insights effectively, impacting their robustness and accuracy. Second, hyperbolic recommender systems lack a systematic approach to balancing exploration (introducing new content) and exploitation (focusing on existing preferences) (Wilson et al., 2021). This limitation hinders their adaptability to users’ evolving interests and reduces the diversity and relevance of recommended items.
Proposed Methods. To address these challenges, we propose HARec, a hyperbolic representation learning framework that facilitates exploration-exploitation through an intuitive tree-structured representation, where depth search enables exploitation and breadth search allows exploration. Our framework introduces two key components: (1) a hierarchical-aware graph-llm alignment mechanism that enhances semantic understanding and reduces noise sensitivity by jointly aligning textual descriptions with user-item collaborative information in hyperbolic space, and (2) a hyperbolic hierarchical tree structure that enables user-adjustable exploration-exploitation trade-offs. This tree structure organizes preferences hierarchically, with higher layers representing broader interests, allowing systematic navigation between exploration and exploitation.
@@
To validate the effectiveness of HARec, we conducted extensive experiments across multiple dimensions. Our model consistently outperformed baselines in both utility and diversity, marking a significant achievement as it is the first model to excel in both aspects simultaneously. To further highlight the benefits of our hyperbolic alignment framework, we compared HARec with feature-empowered baselines, where it demonstrated superior performance. Additionally, we conducted detailed analysis on head and tail items, showing that HARec effectively boosts recommendations for tail items, thus enhancing diversity. An ablation study, along with a decoder module analysis, confirms HARec’s capability to leverage both semantic meaning and collaborative information, providing deeper insights into the advantages of our approach.
Contributions. To summarize, our primary contribution are in three aspects. The first is the proposed HARec, a hyperbolic representation learning framework that effectively balances exploration and exploitation by integrating semantic and collaborative information, supported by theoretical analysis demonstrating the advantages of this integration in hyperbolic space. In addition, we design a novel mechanism enabling personalized control over the exploration-exploitation trade-off through a hierarchical tree structure in hyperbolic space, allowing users to dynamically adjust their recommendation preferences. Moreover, through extensive experiments evaluating both accuracy and diversity metrics, we demonstrate that HARec achieves state-of-the-art performance across both criteria - achieving up to 5.49% improvement in utility metrics and 11.39% increase in diversity metrics, making it the first model to excel in both criteria simultaneously.
Figure 1. Exploration and exploitation in hyperbolic space.
2. Related Work
@@
In this section, we review commonly used graph-based collaborative filtering approaches, large language models (LLMs) for recommendation, and hyperbolic representation learning methods.
@@
Graph Collaborative Filtering. Graph Collaborative Filtering has emerged as a dominant approach in recommender systems, with graph neural networks (GNNs) playing a central role. Unlike traditional collaborative filtering techniques, such as matrix factorization (Koren et al., 2009), GNNs capitalize on the relational structure inherent in user-item interactions, enabling a more nuanced understanding of complex dependencies. Prominent graph-based methods, such as NGCF (Wang et al., 2019), SGL (Wu et al., 2021), and LightGCN (He et al., 2020), have been widely adopted and have consistently achieved state-of-the-art performance, significantly boosting recommendation accuracy. Nevertheless, most of these models are constrained to Euclidean space, which may fall short in capturing the underlying power-law distribution commonly seen in user-item networks.
@@
Large Language Models for Recommendation. Large language models (LLMs) are increasingly used in recommender systems due to their strong capability in understanding textual information. Some approaches leverage LLMs directly for various downstream recommendation tasks (Bao et al., 2023; Li et al., 2023; Lyu et al., 2023; Zheng et al., 2024; Zhang et al., 2023), while others utilize LLMs as an auxiliary tool for collaborative filtering (Ren et al., 2024; Xi et al., 2024; Wei et al., 2024). However, both approaches often require substantial inference time and are computationally inefficient. In our HARec, we use language models solely to enhance data quality, avoiding the need for repeated inference during each recommendation by relying on hyperbolic embeddings for the final output.
@@
3. PRELIMINARIES
Embeddings learned from the above graph collaborative filtering are by default in the Euclidean space, which is unable to capture tree-like structures. To address the problem, we introduce how to derive these embeddings in a hyperbolic space.
@@ 3. PRELIMINARIES
Lorentz Model. An $n$-dimensional Lorentz manifold with negative curvature $-1/\kappa$ $(\kappa>0)$ is described as the Riemannian manifold $(\mathbb{H}^{n}_{\kappa},g_{\mathcal{L}})$. Here, $\mathbb{H}^{n}_{\kappa}=\{x\in\mathbb{R}^{n+1}:\langle x,x\rangle_{\mathcal{L}}=-\kappa,x_{0}>0\}$, $g_{\mathcal{L}}=\eta$ with $\eta=\mathbf{I}_{n}$ except $\eta_{0,0}=-1$, and $\langle\cdot,\cdot\rangle_{\mathcal{L}}$ represents the Lorentzian inner product, which is defined as by
(1) $\langle x,y\rangle_{\mathcal{L}}:=-x_{0}y_{0}+\sum_{i=1}^{n}x_{i}y_{i}.$
@@
(2) $\mathcal{T}_{x}\mathbb{H}_{\kappa}^{n}:=\{v\in\mathbb{R}^{n+1}:\langle v,x\rangle_{\mathcal{L}}=0\}.$
The tangent space is defined to be orthogonal to $x$ with respect to the Lorentzian inner product, therefore making the approximation accurate at the reference point.
@@
(3) $\exp_{x}^{\kappa}(v)=\cosh\left(\frac{\|v\|_{\mathcal{L}}}{\sqrt{\kappa}}\right)x+\sqrt{\kappa}\sinh\left(\frac{\|v\|_{\mathcal{L}}}{\sqrt{\kappa}}\right)\frac{v}{\|v\|_{\mathcal{L}}},$
where $\|v\|_{\mathcal{L}}=\sqrt{\langle v,v\rangle_{\mathcal{L}}}$ denotes the Lorentzian norm of $v$. The logarithmic map $\log_{x}$ is the inverse of the exponential map $\exp_{x}$:
(4) $\log_{x}^{\kappa}(y)=d_{\mathcal{L}}^{\kappa}(x,y)\frac{y+\frac{1}{\kappa}\langle x,y\rangle_{\mathcal{L}}x}{\|y+\frac{1}{\kappa}\langle x,y\rangle_{\mathcal{L}}x\|},$
where $d_{\mathcal{H}}^{\kappa}(x,y)$ represents the distance between $x$ and $y$ in $\mathbb{H}^{n}_{\kappa}$:
(5) $d_{\mathcal{H}}^{\kappa}(x,y)=\sqrt{\kappa}\,\text{arcosh}\left(-\langle x,y\rangle_{\mathcal{L}}/\kappa\right).$
@@
4. METHODOLOGY
In this section, we present a comprehensive overview of our proposed model, HARec. Our approach not only aims to deliver accurate recommendations but also addresses users’ growing demand for more diverse suggestions. Furthermore, by constructing a hierarchical tree from the learned embeddings, we introduce an additional mechanism that enables users to control the degree of exploration in the recommended items. The overall framework is in Figure 2.
Figure 2. The overall architecture of HARec. (i) Hyperbolic Graph Collaborative Filtering: Encodes collaborative information using hyperbolic GNNs; (ii) Hyperbolic Alignment: Aligns semantic and collaborative information within hyperbolic space; (iii) Hyperbolic Hierarchical Structure: Builds hierarchy tree from hyperbolic embeddings.
4.1. Hyperbolic Graph Collaborative Filtering
Similar to Euclidean space, hyperbolic graph collaborative filtering captures high-order dependencies between users and items via a message aggregation mechanism. In our method, we use the Lorentz representation for both users and items.
@@ 4.1. Hyperbolic Graph Collaborative Filtering
(7) $z_{u}^{l}=z_{u}^{l-1}+\sum_{i\in\mathcal{N}_{u}}\frac{1}{|\mathcal{N}_{u}|}z_{i}^{l-1},\quad\quad z_{i}^{l}=z_{i}^{l-1}+\sum_{u\in\mathcal{N}_{i}}\frac{1}{|\mathcal{N}_{i}|}z_{u}^{l-1}$
@@
(9) $h_{u}=\exp_{\mathbf{o}}(z_{u}),\quad\quad h_{i}=\exp_{\mathbf{o}}(z_{i})$
Hyperbolic Prediction Function. Through hyperbolic messaging passing, we’ve captured high-order relational information into hyperbolic embeddings. Hyperbolic distance is now used to define a prediction function, which is used to predict the possibility that a user would interact with an item:
(10) $p(u,i)=\frac{1}{d_{\mathcal{H}}(h_{u},h_{i})}$
Hyperbolic Margin Ranking Loss. Bayesian Personalized Ranking (BPR) loss has been widely applied in recommender systems (Rendle et al., 2012), aiming to pull together user-item pairs with recorded interactions while pushing apart those without. In our model, we adopt the hyperbolic margin ranking loss (Sun et al., 2021), designed to prevent the undesirable collapse seen in Euclidean space by adding an additional margin to the original BPR loss:
(11) $\ell_{margin}(u,i,j)=\max(d_{\mathcal{H}}^{2}(h_{u},h_{i})-d_{\mathcal{H}}^{2}(h_{u},h_{j})+m,0),$
@@
Hyperbolic Informative Mechanism. We further integrate a geometric-aware mechanism (Yang et al., 2022a) to enhance the pull-push process, consisting of two parts: Hyperbolic-aware Margin Learning (HAML) and Hyperbolic Informative Negative Sampling (HINS). Given that distances in hyperbolic space grow exponentially as embeddings deviate from the origin, we aim to assign a larger margin when user-item pairs are near the origin. HAML addresses this issue by dynamically adapting the margin as follows:
(12) $m^{\mathcal{H}}_{ui}$ $=\mathrm{sigmoid}(\delta),$
@@
where $e_{u,0}$ and $e_{i,0}$ represent the zeroth coordinate elements of $e_{u}$ and $e_{i}$, respectively, corresponding to the embeddings norms in the hyperbolic space. HINS is an algorithm that samples negative nodes close to the positive item, yielding informative negative samples for both head and tail items. This is achieved by iteratively calculating the distances between the positive item and all candidate negative samples until the nearest one is identified.
4.2. Semantic Representations Generation
In this subsection, we introduce a paradigm for distilling useful information from raw text descriptions of items and users into dense semantic embeddings, which will be used later in our alignment framework that derives the final embeddings, integrating both collaborative information and semantic insights. Inspired by prior research (Ma et al., 2024; Ren et al., 2024) that leverages the power of large language models (LLMs) to extract useful user/item descriptions from raw, noisy textual data, we adopt a similar framework to generate user/item profiles with rich textual descriptions:
(13) $\mathcal{P}_{i}=\textit{LLMs}(\mathcal{S}_{i},\mathcal{T}_{i})$
where $\mathcal{S}_{i}$ represents system prompts to the LLMs, $\mathcal{T}_{i}$ denotes raw textual information, and $\mathcal{P}_{i}$ is the resulting user/item profile, consisting of sentence-level descriptions.
While we’ve derived textual information for users/items, it cannot be directly fused with the embeddings generated from hyperbolic collaborative filtering. To integrate both types of knowledge, we encode the textual information into embeddings using pretrained text encoder models, such as BERT (Devlin, 2018):
(14) $e_{i}=\textit{Encoder}(\mathcal{P}_{i})$
4.3. Hyperbolic Alignment
Given two sets of embeddings: hyperbolic embeddings with collaborative information and semantic embeddings with textual information, we aim to maximize the mutual information between them, creating unified embeddings that integrate both types of knowledge. To achieve this, we introduce a hyperbolic alignment loss, which mitigates the effects of noisy signals in representation learning and improves recommendations for tail items by aligning information within the same hyperbolic space. A theoretical proof of the advantages of alignment in hyperbolic space over Euclidean space is provided in the next subsection.

### v2

# Breaking Information Cocoons: A Hyperbolic Graph-LLM Framework for Exploration and Exploitation in Recommender Systems

@@ Abstract
Modern recommender systems often create information cocoons, restricting users’ exposure to diverse content. A key challenge lies in balancing content exploration and exploitation while allowing users to adjust their recommendation preferences. Intuitively, this balance can be modeled as a tree-structured representation, where depth search facilitates exploitation and breadth search enables exploration. However, existing approaches face two fundamental limitations: Euclidean methods struggle to capture hierarchical structures, while hyperbolic methods, despite their superior hierarchical modeling, lack semantic understanding of user and item profiles and fail to provide a principled mechanism for balancing exploration and exploitation. To address these challenges, we propose HERec, a hyperbolic graph-LLM framework that effectively balances exploration and exploitation in recommender systems. Our framework introduces two key innovations: (1) a hierarchical-aware graph-LLM mechanism that jointly aligns textual descriptions with user-item collaborative information in hyperbolic space, and (2) a hierarchical representation structure that enables user-adjustable exploration-exploitation trade-offs. Extensive experiments demonstrate that HERec consistently outperforms both Euclidean and hyperbolic baselines, achieving up to 5.49% improvement in utility metrics and 11.39% increase in diversity metrics, effectively mitigating information cocoons.
1 INTRODUCTION
@@
As online platforms grow in size dramatically, users are overwhelmed by an influx of information, creating information cocoons that restrict their exposure to diverse and novel content (Kunaver & Požrl, 2017; Castells et al., 2021).
@@
Traditional recommendation approaches have primarily relied on collaborative filtering and interaction data (He et al., 2020; Koren et al., 2009), where users with similar historical preferences are expected to share future interests (Chen et al., 2021). However, these methods face two critical challenges in improving user experience. First, existing models struggle to capture the underlying hierarchical structure inherent in user-item networks (Zheng et al., 2017; Unger & Tuzhilin, 2020), a critical factor for enhancing performance and personalization. Second, these models fail to effectively balance the exploration-exploitation trade-off. For instance, a jazz enthusiast may want to explore rock music, but current models primarily focus on exploitation—recommending only familiar genres—without adequately supporting exploration (as illustrated in Figure 1).
@@
Hyperbolic space has emerged as a promising approach for modeling hierarchical structures and power-law distributions (Krioukov et al., 2010; Yang et al., 2022b; Sun et al., 2021; Yang et al., 2022a), where its exponential volume growth naturally aligns with hierarchical, scale-free structures (Ganea et al., 2018; Chami et al., 2019; Liu et al., 2019; Zhang et al., 2021; Yang et al., 2023). In hyperbolic space, distances increase exponentially from the origin, allowing popular items or exploratory users to be positioned closer to the origin, while niche items or users with focused interests spread toward the boundaries. This spatial configuration enables hyperbolic graph neural network (GNN) models to perceive hierarchical relationships effectively, enhancing recommendation performance (Yang et al., 2022a; Sun et al., 2021; Yang et al., 2022b).
@@
Despite these advantages, current hyperbolic recommender systems still face key limitations. First, they struggle with a lack of semantic understanding of user and item profiles. While large language models (LLMs) offer a potential solution, their text encoders operate in Euclidean space, making it challenging to integrate LLM-derived semantic features with hyperbolic collaborative signals. This mismatch leaves existing hyperbolic models vulnerable to noise in collaborative information, as they fail to capture semantic insights effectively. Second, hyperbolic recommender systems lack a principled approach to balancing exploration (introducing new content) and exploitation (focusing on existing preferences) (Wilson et al., 2021). This limitation reduces their ability to adapt to the exploration-exploitation trade-off, ultimately diminishing both the utility and diversity of recommendations.
Proposed Methods. To address these challenges, we propose HERec, a hyperbolic graph-LLM framework designed to balance exploration and exploitation in recommender systems. Our approach introduces two key innovations: (1) a hierarchical-aware graph-LLM alignment mechanism, which jointly aligns textual descriptions with user-item collaborative information in hyperbolic space, supported by theoretical gradient analysis demonstrating its advantages, and (2) a novel hierarchical representation structure that enables user-adjustable exploration-exploitation trade-offs. Unlike conventional approaches that require manually defined hierarchies, we propose a hyperparameter-free clustering mechanism theoretically optimized by Dasgupta’s cost (Dasgupta, 2016). This approach automatically discovers hierarchical structures without requiring predefined hyperparameters, allowing for a more adaptive and principled organization of user preferences.
@@
To validate the effectiveness of HERec, we conducted extensive experiments across multiple dimensions. Our model consistently outperformed baselines in both utility and diversity metrics, marking a significant achievement as it is the first model to excel in both aspects simultaneously. For a fair comparison, we enhanced Euclidean baselines with feature information, yet our model still demonstrated superior performance, validating the advantage of our designed alignment in hyperbolic space. Additionally, we conducted a detailed analysis on head and tail items, showing that HERec effectively boosts recommendations for tail items, thus enhancing diversity. An ablation study, along with a hierarchical representation structure analysis and decoder module, confirms HERec’s capability to leverage both semantic meaning and collaborative information, providing deeper insights into the advantages of our approach.
Contributions. Our primary contributions focus on three key aspects: capturing collaborative signals, integrating semantic information, and modeling the underlying hierarchical structure. The proposed graph-LLM alignment framework unifies collaborative and semantic information in hyperbolic space, while a hierarchical representation structure organizes user preferences, allowing for dynamic adjustment between exploration and exploitation. Moreover, through extensive experiments, we demonstrate that HERec achieves state-of-the-art performance across both utility and diversity metrics, making it the first model to excel in both criteria simultaneously. To facilitate reproducibility and further research, we open-source our model implementation at: https://github.com/Martin-qyma/HERec.
Figure 1: Exploration and exploitation in hyperbolic space.
2 Related Work
@@
In this section, we review graph-based collaborative filtering, large language models (LLMs) for recommendation, and hyperbolic representation learning methods.
@@
Graph Collaborative Filtering. Graph Collaborative Filtering has emerged as a dominant approach in recommender systems, with graph neural networks (GNNs) playing a central role. Unlike traditional collaborative filtering techniques, such as matrix factorization (Koren et al., 2009), GNNs capitalize on the relational structure in user-item interactions, enabling a more nuanced understanding of complex dependencies. Prominent graph-based methods, such as NGCF (Wang et al., 2019), SGL (Wu et al., 2021), and LightGCN (He et al., 2020), have been widely adopted and have consistently achieved state-of-the-art performance. Nevertheless, these models are constrained to Euclidean space, which may fall short in capturing the underlying power-law distribution in user-item networks.
@@
Large Language Models for Recommendation. Large language models (LLMs) are increasingly used in recommender systems for their strong capability in understanding textual information. Some approaches leverage LLMs directly for downstream recommendation tasks (Bao et al., 2023; Li et al., 2023; Lyu et al., 2023; Zheng et al., 2024; Zhang et al., 2023), while others utilize LLMs as an auxiliary tool for collaborative filtering (Ren et al., 2024; Xi et al., 2024; Wei et al., 2024). However, these approaches require substantial inference time and are computationally inefficient. In our HERec, we use language models exclusively to improve data quality, eliminating the need for repeated inference during each recommendation by utilizing hyperbolic embeddings for the final output.
@@
3 PRELIMINARIES
Embeddings learned from graph collaborative filtering are by default in the Euclidean space, which is unable to capture tree-like structures. To address the problem, we introduce how to derive these embeddings in a hyperbolic space.
@@ 3 PRELIMINARIES
Lorentz Model. An $n$-dimensional Lorentz manifold with negative curvature $-1/\kappa$ $(\kappa>0)$ is described as the Riemannian manifold $(\mathbb{H}^{n}_{\kappa},g_{\mathcal{L}})$. Here, $\mathbb{H}^{n}_{\kappa}=\{x\in\mathbb{R}^{n+1}:\langle x,x\rangle_{\mathcal{L}}=-\kappa,x_{0}>0\}$, $g_{\mathcal{L}}=\eta$ with $\eta=\mathbf{I}_{n}$ except $\eta_{0,0}=-1$, and $\langle\cdot,\cdot\rangle_{\mathcal{L}}$ represents the Lorentzian inner product, which is defined
$\langle x,y\rangle_{\mathcal{L}}:=-x_{0}y_{0}+\sum_{i=1}^{n}x_{i}y_{i}$ (1)
@@
$\mathcal{T}_{x}\mathbb{H}_{\kappa}^{n}:=\{v\in\mathbb{R}^{n+1}:\langle v,x\rangle_{\mathcal{L}}=0\}$ (2)
@@
$\exp_{x}^{\kappa}(v)=\cosh\left(\frac{\|v\|_{\mathcal{L}}}{\sqrt{\kappa}}\right)x+\sqrt{\kappa}\sinh\left(\frac{\|v\|_{\mathcal{L}}}{\sqrt{\kappa}}\right)\frac{v}{\|v\|_{\mathcal{L}}}$ (3)
where $\|v\|_{\mathcal{L}}=\sqrt{\langle v,v\rangle_{\mathcal{L}}}$ denotes the Lorentzian norm of $v$. The logarithmic map $\log_{x}$ is the inverse of the exponential map $\exp_{x}$, and the distance between $x$ and $y$ in $\mathbb{H}^{n}_{\kappa}$ is:
$\log_{x}^{\kappa}(y)=d_{\mathcal{L}}^{\kappa}(x,y)\frac{y+\frac{1}{\kappa}\langle x,y\rangle_{\mathcal{L}}x}{\|y+\frac{1}{\kappa}\langle x,y\rangle_{\mathcal{L}}x\|}$ (4)
where $d_{\mathcal{H}}^{\kappa}(x,y)$ is the distance between $x$ and $y$ in $\mathbb{H}^{n}_{\kappa}$:
$d_{\mathcal{H}}^{\kappa}(x,y)=\sqrt{\kappa}\,\text{arcosh}\left(-\langle x,y\rangle_{\mathcal{L}}/\kappa\right)$ (5)
@@
4 METHODOLOGY
In this section, we present a comprehensive overview of our proposed model, HERec. Our approach not only aims to deliver accurate recommendations but also addresses users’ growing demand for more diverse suggestions. Furthermore, by constructing a hierarchical tree from the learned embeddings, we introduce an additional mechanism that enables users to control the degree of exploration in the recommended items. The overall framework is in Figure 2.
Figure 2: The overall architecture of HERec. (i) Hyperbolic Graph Collaborative Filtering: Encodes collaborative information using hyperbolic GNNs; (ii) Hyperbolic Alignment: Aligns semantic and collaborative information within hyperbolic space; (iii) Hierarchical Representation Structure: Builds hierarchy structure from hyperbolic embeddings.
4.1 Hyperbolic Graph Collaborative Filtering
Similar to Euclidean space, hyperbolic graph collaborative filtering captures user-item dependencies via message passing mechanism. In our method, we use the Lorentz representation for both users and items.
@@ 4.1 Hyperbolic Graph Collaborative Filtering
$z_{u}^{l}=z_{u}^{l-1}+\sum_{i\in\mathcal{N}_{u}}\frac{1}{|\mathcal{N}_{u}|}z_{i}^{l-1},\ z_{i}^{l}=z_{i}^{l-1}+\sum_{u\in\mathcal{N}_{i}}\frac{1}{|\mathcal{N}_{i}|}z_{u}^{l-1}$ (7)
@@
$h_{u}=\exp_{\mathbf{o}}(z_{u}),\quad\quad h_{i}=\exp_{\mathbf{o}}(z_{i})$ (9)
Hyperbolic Prediction Function. Through hyperbolic messaging passing, we’ve captured high-order relational information into hyperbolic embeddings. Hyperbolic distance is now used to define a prediction function, which predicts the possibility that a user would interact with an item:
$p(u,i)=\frac{1}{d_{\mathcal{H}}(h_{u},h_{i})}$ (10)
Hyperbolic Margin Ranking and Informative Mechanism. To address the limitations of Euclidean-based approaches, we adopt the hyperbolic margin ranking loss (Sun et al., 2021), an extension of Bayesian Personalized Ranking (BPR) (Rendle et al., 2012). This loss pulls user-item pairs with observed interactions closer while pushing negative pairs apart, preventing the collapse seen in Euclidean space by incorporating a margin:
$\ell_{m}(u,i,j)=\max(d_{\mathcal{H}}^{2}(h_{u},h_{i})-d_{\mathcal{H}}^{2}(h_{u},h_{j})+m,0)$ (11)
@@
To further refine this pull-push process, we integrate a geometric-aware mechanism (Yang et al., 2022a) combining Hyperbolic-aware Margin Learning (HAML) and Hyperbolic Informative Negative Sampling (HINS). HAML dynamically adjusts the margin based on hyperbolic geometry, assigning a larger margin to pairs near the origin:
$m^{\mathcal{H}}_{ui}$ $=\mathrm{sigmoid}(\delta)$ (12)
@@
where $e_{u,0}$ and $e_{i,0}$ are the zeroth coordinate elements of $e_{u}$ and $e_{i}$, respectively, corresponding to the embeddings norms in the hyperbolic space. HINS enhances negative sampling by prioritizing candidates close to the positive item, iteratively identifying the most informative negative samples. This method leverages hyperbolic geometry to effectively model head and tail items, improving utility performance in recommendations.
4.2 Semantic Representations Generation
In this subsection, we introduce a paradigm for distilling useful information from raw text descriptions of items and users into dense semantic embeddings, which will be used in our alignment framework to derive the final representation. Inspired by prior research (Ma et al., 2024; Ren et al., 2024) that leverages the power of large language models (LLMs) to extract useful user/item descriptions from raw, noisy textual data, we adopt a similar framework to generate user/item profiles with rich textual descriptions: $\mathcal{P}_{i}=\textit{LLMs}(\mathcal{S}_{i},\mathcal{T}_{i})$, where $\mathcal{S}_{i}$ represents system prompts to the LLMs, $\mathcal{T}_{i}$ denotes raw textual information, and $\mathcal{P}_{i}$ is the resulting user/item profile, consisting of sentence-level descriptions.
While we’ve derived textual information for users/items, it cannot be directly fused with the embeddings generated from hyperbolic collaborative filtering due to discrepancies in embedding geometry. To integrate both types of knowledge, we encode the textual information in embeddings using pre-trained text encoder models: $e_{i}=\textit{Encoder}(\mathcal{P}_{i})$. While LLM embeddings could be an alternative, they often introduce higher computational overhead and may not be optimized for alignment with hyperbolic representations, making encoders a more practical choice.
4.3 Hyperbolic Alignment
We propose maximizing mutual information between hyperbolic embeddings (collaborative information) and semantic embeddings (textual information) to create unified embeddings that integrate both knowledge types. To achieve this, we introduce a hyperbolic alignment loss, which reduces noise in representations and improves tail item recommendations by aligning information in the same hyperbolic space. A theoretical advantage of hyperbolic alignment over Euclidean alignment is discussed in the next subsection.